# LiLIS: Enhancing Big Spatial Data Processing with Lightweight Distributed Learned Index

### v1

@@ 5.1.4 Research Questions
5.2 RQ1: Overall Performance
To answer RQ1, we compare LiLIS-K with all other alternative big spatial analytics for range queries, point queries, and $k$NN under the default settings. Because Quad-tree index in Sedona does not support $k$NN, we only report Sedona-RQ and Sedona-RK for $k$NN queries. As for join queries, since the spatial join predicate used in LiLIS is not found in Sedona, we only compare LiLIS-K with Spark. The evaluation results are illustrated in Figure 4. As we can see, LiLIS performs consistently the best compared with other implementations. Since the y-axis in Figure 4 uses a logarithmic scale, the actual performance improvement is significantly greater than the observed values suggest. For example, LiLIS-K takes about 472 millisecond for range queries, while Sedona-RQ takes 521282 millisecond. Interestingly, we can also observe that Sedona variants with indices are not better than Sedona-N and Spark under our settings, and this phenomenon is caused by the underlying data distribution as well as its Filter and Refine model. The authors of Sedona [7] also verify this finding.
@@ 5.2 RQ1: Overall Performance
Takeaway 1 LiLIS is the fastest, and it outperforms all competitors by 2–3 orders of magnitude.
5.3 RQ2: Varying Partitioners
An appropriate partitioner is essential for the best performance. Since the name of a LiLIS variant can indicate the chosen paritioner, we denote the paritioner as its implementation directly in Table 3 for the sake of simplicity. We highlight the best results in light green and the second-best results in yellow. As we can see, both Quad-tree and KD-tree are ideal paritioners for all queries, and R-tree paritioner can be seen as the best choice except for join queries. This difference highlights a fundamental tradeoff: while R-tree optimizes for individual object retrieval, the join operation’s combinatorial nature exposes weaknesses in its structure that Quad-tree and KD-tree handle more gracefully due to their more uniform division strategies.
Table 3: Varying Partitioners (ms)
Paritioner Query Type Point Query Range Query $k$NN Query Join Query
@@ 5.4 RQ3: Varying Datasets
The results in Figure 5 also indicate that $k$NN is more time-consuming than both point and range query. Therefore, we further take it as the running example to compare LiLIS with its competitors among various datasets. In Table 4, we highlight the best results in bold, demonstrating that LiLIS consistently outperforms other methods across all datasets.
Table 4: Varying Datasets for $kNN$ Queries (ms)
@@
5.5 RQ4: Subtle Factors for Range and $k$NN Queries
@@
Next, we further conduct experiments over all variants of LiLIS while varying $k$ from 1 to 100 for $k$NN queries. According to the results in Figure 7, we can conclude that $k$NN queries is relatively stable when varying $k$, and this is consistent with the analysis presented in Section 4.3. This stability can be attributed to the query processing mechanism: regardless of how $k$ changes, the number of partitions that need to be accessed remains relatively small, as the query radius determination primarily depends on data distribution rather than the $k$ value itself. Additionally, we can also notice that R-tree partitioner consistently outperforms other partitioning strategies across different $k$ values. This superior performance can be explained by the R-tree’s effective spatial indexing capabilities, which enable more efficient pruning of irrelevant partitions during query processing, thereby reducing the overall computational cost.
@@ 5.6 RQ5: Index Cost
To demonstrate the superior efficiency of LiLIS in index construction, we conduct comparative experiments measuring the build time for all index variants in both LiLIS and Sedona. As illustrated in Figure 8, LiLIS consistently achieves faster index construction times compared to traditional spatial indices. For instance, building a LiLIS-K index requires approximately 50 seconds, whereas constructing a Sedona-RQ index takes about 100 seconds, representing a 1.5-2.0$\times$ speed-up. Notably, while the performance advantage is significant, the gap in construction time is less pronounced than the query time improvements we observed above. This observation aligns with our theoretical analysis presented in Section 3.2.

### v3

@@ V-A4 Research Questions
V-B RQ1: Overall Performance
To answer RQ1, we compare LiLIS-K with all other alternative big spatial analytics for range queries, point queries, and $k$NN under the default settings. Because Quad-tree index in Sedona does not support $k$NN, we only report Sedona-RQ and Sedona-RK for $k$NN queries. As for join queries, since the spatial join predicate used in LiLIS is not found in Sedona, we only compare LiLIS-K with Apache Spark. The evaluation results are illustrated in Figure 4. As we can see, LiLIS performs consistently the best compared with other implementations. Since the y-axis in Figure 4 uses a logarithmic scale, the actual performance improvement is significantly greater than the observed values suggest. For example, LiLIS-K takes about 472 millisecond for range queries, while Sedona-RQ takes 521282 millisecond. Interestingly, we can also observe that Sedona variants with indices are not better than Sedona-N and Spark under our settings, and this phenomenon is caused by the underlying data distribution as well as its Filter and Refine model. The authors of Sedona [9] also verify this finding.
@@ V-B RQ1: Overall Performance
Takeaway 1 LiLIS is the fastest for spatial points, and it generally outperforms competitors by 2–3 orders of magnitude.
As for RQ1, We also conducted the experiments in terms of throughput (jobs per minute) of LiLIS and its competitors of point and range queries. To be specific, we followed the similar settings in Simba [8], and issued 100 queries on the driver program using a thread-pool with 8 threads. The results are shown in Figure 5. It can be observed that LiLIS-K consistently achieves higher throughput, but the advantage is not as obvious as that in query time, and this is mainly because the throughput is influenced by additional factors such as I/O overhead, CPU cores, and memory bandwidth that are not solely dependent on query execution efficiency.
Figure 5: The throughput (jobs per minute) of point and range queries under default settings.
V-C RQ2: Varying Partitioners
An appropriate partitioner is essential for the best performance. Since the name of a LiLIS variant can indicate the chosen paritioner, we denote the paritioner as its implementation directly in Table III for the sake of simplicity. We highlight the best results in light green and the second-best results in yellow. Our analysis reveals that both Quad-tree and KD-tree serve as highly effective partitioning strategies across a broad range of query types, while the R-tree partitioner emerges as the optimal choice for most queries, with the notable exception of join queries. This distinction underscores a critical tradeoff in spatial indexing: the R-tree’s optimization for efficient individual object retrieval is less effective for join operations, where its structural characteristics are less suited to the combinatorial demands of such queries. In contrast, the more uniform spatial division strategies employed by Quad-tree and KD-tree enable them to handle join queries more adeptly, highlighting their robustness in scenarios requiring complex relational operations.
TABLE III: Varying Partitioners (ms)
Method Query Point Query Range Query $k$NN Query Join Query
@@ V-D RQ3: Varying Datasets
The results in Figure 6 also indicate that $k$NN is more time-consuming than both point and range query. Therefore, we further take it as the running example to compare LiLIS with its competitors among various datasets. In Table IV, we highlight the best results in bold, demonstrating that LiLIS consistently outperforms other methods across all datasets.
TABLE IV: Varying Datasets for $kNN$ Queries (ms)
@@
V-E RQ4: Subtle Factors for Range and $k$NN Queries
@@
Next, we further conduct experiments over all variants of LiLIS while varying $k$ from 1 to 100 for $k$NN queries. According to the results in Figure 8, we can conclude that $k$NN queries is relatively stable when varying $k$, and this is consistent with the analysis presented in Section IV-C. This stability can be attributed to the query processing mechanism: regardless of how $k$ changes, the number of partitions that need to be accessed remains relatively small, as the query radius determination primarily depends on data distribution rather than the $k$ value itself. Additionally, we can also notice that R-tree partitioner consistently outperforms other partitioning strategies across different $k$ values. This superior performance can be explained by the R-tree’s effective spatial indexing capabilities, which enable more efficient pruning of irrelevant partitions during query processing, thereby reducing the overall computational cost.
@@ V-F RQ5: Index Cost
To demonstrate the superior efficiency of LiLIS in index construction, we conduct comparative experiments measuring the build time for all index variants in both LiLIS and Sedona. As illustrated in Figure 9, LiLIS consistently achieves faster index construction times compared to traditional spatial indices. For instance, building a LiLIS-K index requires approximately 50 seconds, whereas constructing a Sedona-RQ index takes about 100 seconds, representing a 1.5-2.0$\times$ speed-up. Notably, while the performance advantage is significant, the gap in construction time is less pronounced than the query time improvements we observed above. This observation aligns with our theoretical analysis presented in Section III-B.